OpenStudy (anonymous):

ok I am confused on this problem: p + 9/p = -10

OpenStudy (ybarrap):

p = -9 p = -1

OpenStudy (ybarrap):

by inspection

OpenStudy (anonymous):

(p^2+9)/p = -10 p^2+9=-10p=0 p^2+10p+9=0 p^2+p+9p+9=0 p(p+1)+9(p+1)=0 (p+1)(p+9)=0 p+1=0 so p=-1 p+9= s0 p=-9

OpenStudy (anonymous):

thank you for all your help

OpenStudy (anonymous):

its ok...:P